# RoboCar: A Rapidly Deployable Open-Source Platform for Autonomous Driving Research

### v1

@@ IV-E1 Simulation
RoboCar implements a simulation module designed to help with the development by offering a virtual test ground for the various algorithms and systems involved. The simulation module contains a component emulating an ego vehicle and components simulating virtual object agents. The vehicle simulation component is based on the kinematic bicycle model[37] and on a simple throttle and steering emulation. The development of a more advanced vehicle model is part of the future work. The virtual object agents can be implemented by the user to tailor specific behaviour and create adequate driving scenarios.
@@ IV-F Main Data Flows
Fig. 6 depicts an overview of the main data flows in RoboCar. The images and the point clouds are supplied by the sensors drivers to the perception, which in turn provides the motion planner with spatially detected objects. The localization is fed with raw positions coming from the GNSS driver and then produces a referenced position in the map frame. This position is used by the map module to select relevant waypoints according to the routing. In addition to the objects, the position and the waypoints are fed to the motion planner, which computes an optimal and safe trajectory. The control receives both the trajectory and the position and computes an optimal actuation commands to properly follow the trajectory. The actuation commands are then sent to the vehicle module which writes the CAN messages to the CAN driver before repeating the entire progress.
@@ V-B High-Definition Mapping
The high-definition (HD) map is a critical component of autonomous driving systems, providing centimetric and lane-level information needed for localization, perception, scene understanding and motion planning [25]. HD maps are composed of several layers, each plays one or more of the above-mentioned roles. The base map layer is usually in the form of a point cloud representing the 3D details of the environment and mainly used to improve localization by matching with the LiDAR point cloud. For RoboCar, we created a base map for a test route of about 3km in Kirchberg (Fig. 7.a) using an adapted version of the FAST-LIO algorithm that robustly fuses the LiDAR point cloud of the Ouster OS1 and INS data of the Trimble BX922 using a tightly-coupled iterated Kalman filter [38]. Further, this base layer is georeferenced in QGIS [^2] https://qgis.org/, an open-source geographical information system (GIS) software, where the geometric, semantic and topological layers are created (Fig. 7.b). Finally, all layers are superimposed altogether for use in the ADS (Fig. 7.c).
@@ V-E Experimental setup
In the context of this paper, a dedicated road experiment was conducted on the same route previously described to evaluate the latest capabilities of RoboCar. The speed limit $v_{max}$ was set to 40 km/h and leaving the center lane was not allowed, meaning that the vehicle would have to stop on every blocking obstacle encountered. The traffic light perception relied on the Intel RealSense camera placed inside the cabin on the windshield. The parameters of the RSS used for the experiment are summarized on Table I.
@@
TABLE I: Parameters used for the public road experiment with $v_{max}$ as the speed limit throughout the path. The $\beta_{min}$ is voluntarily defensive for greater safety during the tests.
@@ V-F Results
The experimental vehicle running RoboCar successfully completed the Kirchberg loop while respecting the speed limits and traffic lights. A safe distance from a leading vehicle was also properly maintained requiring no manual intervention. However, there is room for improvement in the smoothness of the car-following motion, as the vehicle tended to oscillate around the safe distance. On one occasion, the intervention of the safety driver was required to correct a slight deviation from the path. The traffic lights were on average detected by the system at a distance of $44m$, which proved sufficient to smoothly stop the vehicle once the signal turned red. Upon investigation, the limiting factors of the detection distance seemed to be the map module, which would not always load the upcoming traffic lights in a timely manner as well as the curvature of the road combined with obstacle obstructing the sight. The perception system in itself was able to detect traffic lights from a distance of up to $70m$. Fig. 9 illustrates a snippet of the results along a section of the Kirchberg route used for the experiments. The speed of the vehicle and an estimation of the center lane deviation are reported alongside snapshots of the front camera. Snapshot 1 and 2 show the vehicle at a cruising phase properly detecting a green traffic light at a distance and observing the speed limit. Some oscillations around the center lane can be observed, however, their amplitude and period make them barely perceptible for a passenger onboard. Snapshot 3 illustrates the vehicle entering the roundabout having significantly reduced its speed. It can be observed throughout the roundabout that the center lane deviation can get more significant at about $0.75m$. Snapshot 4 is taken right before one manual intervention that was required at around $t=69s$ to avoid that the car would drift too far away from the center lane risking to graze the sidewalk. This center lane deviation could be the result of inaccuracies produced by the motion planner and/or the control module. The autonomous driving was re-engaged a few seconds later around $t=72$. Snapshot 5 shows the car fully stopped at a red light while snapshot 6 is taken after the green light with the vehicle being on a smooth acceleration path to another cruising phase. The right turn following the traffic light on snapshot 5 being quite tight, it can again be observed that the deviation was more significant. Overall, the speed transitions were smooth and allowed for a comfortable driving.
@@ VI-A Platform scope
The experiments demonstrated the ability of RoboCar and its hardware platform to enable basic self-driving operations on public roads, however, some limitations remain requiring additional work to reach full autonomy reliably. The hardware platform itself has limitations and also constraints further development due to a lack of sufficient computing power and limited sensors among other things. In this context, RoboCar is not positioned as full-blown replacement of existing autonomous driving solutions targeting production deployments but rather as a foundational framework allowing researchers to develop and test new techniques and algorithms related to connected and autonomous mobility. The relatively small footprint of RoboCar, the modularity and the ROS2 compatibility allow for a quick and easy prototyping while remaining accessible on a broad range of inexpensive hardware.

### v2

@@ IV-E1 Simulation
RoboCar implements a simulation module designed to help with the development by offering a virtual test ground for the various algorithms and systems involved. The simulation module contains a component emulating an ego vehicle and components simulating virtual object agents. The vehicle simulation component is based on the kinematic bicycle model[23] and on a simple throttle and steering emulation. The development of a more advanced vehicle model is part of future work. Virtual object agents can be implemented by the user to tailor specific behaviors and create adequate driving scenarios.
@@ IV-F Main data flows
Fig. 7 depicts an overview of the main data flows in RoboCar. The images and the point clouds are supplied by the sensors drivers to the perception, which in turn provides the motion planner with spatially detected objects. The localization is fed with raw positions coming from the GNSS driver and then produces a referenced position in the map frame. This position is used by the map module to select relevant waypoints according to the routing. In addition to the objects, the position and the waypoints are fed to the motion planner, which computes an optimal and safe trajectory. The control receives both the trajectory and the position and computes optimal actuation commands to properly follow the trajectory. The actuation commands are then sent to the vehicle module which writes the CAN messages to the CAN driver before repeating the entire process.
@@ V-B High-definition mapping
The high-definition (HD) map is a critical component of autonomous driving systems, providing centimetric and lane-level information needed for localization, perception, scene understanding and motion planning [11].
HD maps are composed of several layers, each plays one or more of the above-mentioned roles. The base map layer is usually in the form of a point cloud representing the 3D details of the environment. The geometric layer provides a vectorized lane-level representation of road geometry using basic geometric primitives (points, line segments, and polygons). While the semantic layer assigns semantic features to these primitives, the road connectivity layer defines how these geometric primitives are connected. In the current release of RoboCar, we provide two examples of HD maps that we have created for our public road testing. The first is a public road track of about 3km in Kirchberg (Fig. 8) while the second is a public road track of about 2km in Belval (Fig. 4). In RoboCar, we create the base map layer (Fig. 8.a) using an adapted version of the FAST-LIO mapping algorithm [36] that robustly fuses the LiDAR point cloud of the Ouster OS1 and INS data of the Trimble BX922 using a tightly-coupled iterated Kalman filter. The original FAST-LIO fuses the LiDAR point cloud and IMU to create a point cloud map while simultaneously estimating vehicle odometry and positional data. We take advantage of the high-precision of INS data of the Trimble BX922 positional data as an additional input in the estimation of the map to avoid the well-known loop closure problem. This mapping process generates a point cloud base map in Cartesian coordinates which is then converted to geodesic coordinates via georeferencing. The base map is imported into the geographical information system (GIS) software QGIS [^2] https://qgis.org. The geometric, semantic and road connectivity layers are then created. First, geometric primitives of the geometric layer are manually crafted by matching and identifying features in the point cloud map’s intensity and colors. Next, these elements are traced to assign semantic attributes and define their connections. Finally, the vectorized layers (Fig. 8.b) are exported to a GeoJson format for easy use in RoboCar. All layers are superimposed altogether in Fig. 8.c.
@@ V-E Experimental setup
In the context of this paper, a dedicated road experiment was conducted on the same route previously described to evaluate the latest capabilities of RoboCar. The speed limit $v_{max}$ was set to 40 km/h and leaving the center lane was not allowed, meaning that the vehicle would have to stop on every blocking obstacle encountered. The traffic light perception relied on the Intel RealSense camera placed inside the cabin to validate the system in its simplest and most cost-effective configuration. The information on which traffic light to follow along the path was known in advance and encoded in the map so that the motion planner can select the proper traffic light from the perception. The parameters of the RSS used for the experiment are summarized on Table I.
TABLE I: Public road experiment parameters.
@@ V-F Results
The experimental vehicle running RoboCar successfully completed the Kirchberg loop while respecting the speed limits and traffic lights. A safe distance from a leading vehicle was also properly maintained, requiring no manual intervention. However, there is room for improvement in the smoothness of the car-following motion, as the vehicle tended to oscillate around the safe distance. On one occasion, the intervention of the safety driver was required to correct a slight deviation from the path. The traffic lights were on average detected by the system at a distance of $44m$, which proved sufficient to smoothly stop the vehicle once the signal turned red. Upon investigation, the limiting factors of the detection distance seemed to be the map module, which would not always load the upcoming traffic lights in a timely manner as well as the curvature of the road combined with obstacle obstructing the sight. The perception system in itself was able to detect traffic lights from a distance of up to $70m$. Fig. 10 illustrates a snippet of the results along a section of the Kirchberg route used for the experiments. The speed of the vehicle and an estimate of the center lane deviation are reported alongside snapshots of the front camera. Snapshots 1 and 2 show the vehicle in a cruising phase properly detecting a green traffic light at a distance and observing the speed limit. Some oscillations around the center lane can be observed, however, their amplitude and period make them barely perceptible for a passenger onboard. Snapshot 3 illustrates the vehicle entering the roundabout having significantly reduced its speed. It can be observed throughout the roundabout that the center lane deviation can get more significant at about $0.75m$. Snapshot 4 is taken right before one manual intervention that was required at around $t=69s$ to avoid the car drifting too far away from the center lane, risking grazing the sidewalk. This center lane deviation could be the result of inaccuracies produced by the motion planner and/or the control module. The autonomous driving was re-engaged a few seconds later around $t=72$. Snapshot 5 shows the car fully stopped at a red light while snapshot 6 is taken after the green light with the vehicle being on a smooth acceleration path to another cruising phase. The right turn following the traffic light on snapshot 5 being quite tight, it can again be observed that the deviation was more significant. Overall, the speed transitions were smooth and allowed for a comfortable driving.
@@ VI-A Platform scope
The experiments demonstrated the ability of RoboCar and its hardware platform to enable basic self-driving operations on public roads, however, some limitations remain requiring additional work to reach full autonomy reliably. The hardware platform itself has limitations and also constraints further development due to a lack of sufficient computing power and limited sensors among other things. In this context, RoboCar is not positioned as full-blown replacement of existing autonomous driving solutions targeting production deployments but rather as a foundational framework allowing researchers to develop and test new techniques and algorithms related to connected and autonomous mobility. Compared to more advanced and complex platforms such as Apollo and Autoware, RoboCar is a lightweight, easy-to-deploy autonomous driving stack designed for accessibility and scalability. While Apollo and Autoware offer extensive features, they come with significant complexity, numerous external dependencies, and are not well-suited for the European regulatory and operational landscape. RoboCar, in contrast, has a low footprint and can seamlessly scale from a small RC car to a full-blown autonomous vehicle. Its hardware requirements are minimal, making it compatible with a wide range of inexpensive computing platforms. Unlike Apollo, RoboCar is built on ROS2, enhancing modularity and interoperability with other projects. By focusing on essential functionality without unnecessary overhead, RoboCar offers a practical, efficient, and accessible solution for autonomous driving development.
@@ VI-B Limitations and future works
While RoboCar is designed to be lightweight and cost-effective, scalability and complexity are not limitations but opportunities for growth. Our primary goal is to foster engagement within the European research community, encouraging collaboration through open-source contributions, research partnerships, and potentially EU-funded projects. By providing a solid, easy-to-use foundation, we aim to empower researchers and developers to extend RoboCar for more complex use cases, from urban autonomy to large-scale fleet deployments. Rather than building a monolithic system upfront, we believe in a modular, community-driven evolution, ensuring that RoboCar remains adaptable, interoperable, and well-suited for the unique challenges of autonomous mobility in Europe.